please put into slope intercept form:
X<= 2y +8
2y + 3x <= 4
2y <= 3x +4
Thanks
BB

Answer by tutorcecilia(2152)

X<= 2y +8 [solve for the y-term]
x-2y<=2y-2y+8
x-2y<=8
x-x-2y<=-x+8
-2y<=-x+8
-2y/-2=>-x/-2+8/-2 [flip the inequality sign when multiplying or dividing by a negative number]
.
y => 1/2x-4
.
2y + 3x <= 4 [solve for the y-term]
2y+3x-3x<=-3x+4
2y<=-3x+4
2y/2=-3x/2+4/2
y= -3/2x + 2
.
2y <= 3x +4 [solve for the y-term]
2y/2 <+3x/2+4/2
y<=3/2x+2